OpenStudy (anonymous):

M+4over 3=4m-5 over7

OpenStudy (anonymous):

\[\frac{m+4}{3} = \frac{4m-5}{7}\]Cross multiplication: 7(m+4) = 3(4m-5) 7m + 28 = 12m - 15 ^solve this

OpenStudy (anonymous):